OpenStudy (anonymous):

Change of Variables:Polar Coordinates Evaluate the iterated integral by converting to polar coordinates. sin(x^2+y^2)dydx

OpenStudy (anonymous):

I want to make sure I'll repost my equation integrals correctly, hold on I'll repost my equation

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\int\limits_{0}^{\sqrt{4-x^2}}\sin \sqrt{x^2+y^2}dydx\] This is the problem, I'll post up my limits that I got

OpenStudy (anonymous):

 The limits are this: r= 0,\[\pi\] \[\theta\]= 0,2